OpenStudy (anonymous):

Who can answer this? Medals given:

HanAkoSolo (jamierox4ev3r):

answer what?

OpenStudy (anonymous):



HanAkoSolo (jamierox4ev3r):

wouldn't that simply equal 8x? if you used the reciprocal?

OpenStudy (anonymous):

Can you show me how you got that?

zepdrix (zepdrix):

\[\large \frac{x}{3}=\frac{8}{y}\]Multiply both sides by y,\[\large y\frac{x}{3}=\frac{8}{\cancel y}\cancel y\]The y's on the right will cancel out, giving us,\[\large \frac{xy}{3}=8\]Divide both sides by x,\[\large \frac{\cancel xy}{\cancel x3}=\frac{8}{x}\]Leaving us with,\[\large \frac{y}{3}=\frac{8}{x}\] We basically just traded places with the y and x :) Lemme know if any of that was too confusing.

OpenStudy (anonymous):

Thanks!